OpenStudy (anonymous):

Show that 13^n = 6^n (mod 7) for all n in Z+ using mathematical induction.

OpenStudy (anonymous):

Trying to solve it with P(n+1) = 13^(n+1) = 6^(n+1) (mod 7) and not with the (x^(n+1) - y^(n+1)) way

OpenStudy (anonymous):

Summation